[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 25, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed May 25 11:29:50 UTC 2016


Here is today's summary of NHC computer operations:
--- mello

----------

TAFB:
TAFB noted an isolated problem sending the GFE grids to NAWIPS Monday.  
Networking issues with the comp server 03 caused the problem. NCO is  
working to ensure the problem does not occur again.
HSU:
TSB fixed the problem with the UKMET data flow. Consequently, it also fixed  
the GUNC model (The GUNC still lives !).
Other items:
The Himawari data is flowing again.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160525/55d265ef/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list